OpenStudy (anonymous):

In triangle ABC, a=4, b=7, c=9. Find the measure of angle C to the nearest tenth of a degree. My progress: 9^2=7^2+4^2-2(9)(7)cosC 81=49+16-126cosC 81=65 - 126cosC ? I am unsure how to evaluate it.

OpenStudy (thomas5267):

\[ \begin{align*} 81&=65-126\cos C \\ 16&=-126\cos C \\ \frac{-8}{63}&=\cos C \end{align*} \]

OpenStudy (anonymous):

-0.1269?

OpenStudy (anonymous):

Oh 2nd cos button?

OpenStudy (thomas5267):

Use the inverse cosine function.

OpenStudy (thomas5267):

What is your answer?